Eight years from now, Clifford will be twice his age was 6years ago. What is Clifford's present age?
Answer by ankor@dixie-net.com(22740)

eight years from now, Clifford will be twice his age was 6years ago.
What is Clifford's present age?
:
a = his present age
a + 8 = 2(a-6)
a + 8 = 2a - 12
8 + 12 = 2a - a
20 = a
C is 20 now
:
:
Check
20 + 8 = 2(20-6)
28 = 2(14)